OpenStudy (anonymous):

Determine the slope of the secant line through the given x values. f(x)=5x^4-9x^2+2x-7, x=4, x=11

OpenStudy (anonymous):

f'(x) for each x gives u the slope on that point so u have : 20x^3-18x+2 so put the x in the equation : 20(4)^3-18(4)+2 and do the same for x=11

OpenStudy (anonymous):

thanks dear , hope u be successful in the exams .

OpenStudy (paxpolaris):

find f(4) & f(11)

OpenStudy (anonymous):

thanks. Also, how would you find an equation of the secant line through the given values of x?

OpenStudy (paxpolaris):

find the line that passes through (4,f(4)) & (11,f(11))

OpenStudy (anonymous):

as i remember u should inverse the slope and multiple it with -1 and then with this new slope put the point in this order ; y=mx+a u should put the new slope instead m and u will find a by putting the point in equation .

OpenStudy (anonymous):

Thanks for your help guys :)

OpenStudy (anonymous):

ur welcome dear

OpenStudy (paxpolaris):

the secant passes through: \(\Large \left( 4,\ 1137 \right)\) & \(\Large \left( 11,\ 72131 \right)\) so the equation is \[\Large y= 10142x-39431\] http://www.wolframalpha.com/input/?i=y%3D5x^4-9x^2%2B2x-7%2C+y%3D+10142x-39431%2C+%28x+%3D+4+to+11%29 http://www.wolframalpha.com/input/?i=real+solutions+y%3D5x^4-9x^2%2B2x-7%2C+y%3D+10142x-39431